OpenStudy (anonymous):

My textbook is saying... Observe that F(x)=2^(-x)=f(-x), thus, the graph of F is a reflection (in the y-axis) of the graph of f. How is this possible? Aren't they the same graph? Is there a difference between the symbols F(x) and f(x) that I don't know about? Thanks for any help! Will give medal, of course :)

OpenStudy (anonymous):

\[F(x)=2^{-x}=f ^{-x}\]

OpenStudy (xapproachesinfinity):

Did your textbook say that f(x)=2^x

OpenStudy (anonymous):

@inkyvoyd

OpenStudy (anonymous):

No, just F(x)=2^(-x). F(x) is the same thing as f(x), right?

OpenStudy (xapproachesinfinity):

No there not the same. Could you take a snapshot to your textbook page

OpenStudy (anonymous):

Yeah, I'll try. It will take me a second.

OpenStudy (xapproachesinfinity):

i take it that f(x)=2^x when you do f(-x)=2^(-x)=F(x) and i F(x) is really a reflection of f along the y axis

OpenStudy (anonymous):

Sorry that took a while; my computer's lagging behind

OpenStudy (anonymous):

It's the top paragraph.

OpenStudy (anonymous):

Okay, I see what you said in your last reply. In example 2, is says that f(x)=2^x

OpenStudy (xapproachesinfinity):

i see, the thing is, they have talked about f(x) in the previous example with g(x) so f(x)=2^x, g(x)=4^x check your textbook for this haha

OpenStudy (xapproachesinfinity):

well if you want to understand this try graphing 2^x and 2^-x and see how the graph would look like

OpenStudy (xapproachesinfinity):



OpenStudy (xapproachesinfinity):

it should look this way!